OpenStudy (anonymous):

A cone has a diameter of 6 centimeters and a height that is 3 times the diameter. Using 3.14 for pi, which of the following can be used to calculate the volume of the cone? one over three(3.14)(3cm)2(18cm) one over three(3.14)(6cm)2(18cm) one over three(3.14)(18cm)2(3cm) one over three(3.14)(18cm)2(6cm)

OpenStudy (anonymous):

@geekfromthefutur

OpenStudy (anonymous):

@Jack1

OpenStudy (jack1):

From Question: "A cone has a diameter of 6 centimeters" therefore Radius = half of diameter radius = 3 cm

OpenStudy (jack1):

From question: "and a height that is 3 times the diameter" h = 3D D = 6 therefore h = 18

OpenStudy (jack1):

so if Volume of a Cone = pi x r^2 x h x 1/3 \(\Large V = \pi \times r^2 \times h \times \frac13\)

OpenStudy (jack1):

so which answer do you thing it is @Moo_Moo17 ?

OpenStudy (anonymous):

i think its A (:

OpenStudy (jack1):

yep, perfect!

OpenStudy (anonymous):

oh ok i only have 3 more question ^-^ would it be ok if u could help me ??